Felicia can sort a paper in 5 hrs and together Tarance can sort it in 2 hrs so what extent would it take Tarance to sort the same paper alone? Doing Rational Equations but confuse on how to start this one.
Answer by stanbon(75887)

Felicia can sort a paper in 5 hrs and together Tarance can sort it in 2 hrs so what extent would it take Tarance to sort the same paper alone? Doing Rational Equations but confuse on how to start this one.
-----
Relicia DATA:
time = 5 hrs/job ; rate = 1/5 job/hr
--------------
Terance DATA:
time = 2 hrs/job ; rate = 1/2 job/hr
------------
Together DATA:
time = 2 hrs/job ; rate = 1/x job/hr
---------------
Equation:
rate + rate = together rate
1/5 + 1/2 = 1/x
------
2x + 5x = 10
7x = 10
----
x = 10/7 hrs = 1.4286 hours = 1hr 25min 42sec (time to do the job together)
------------
Cheers,
Stan H.
------------